# Positive operator-valued noncommutative polynomials are squares

### v2

@@ 1.1. Notation
Let $\mathcal{H}$ be a fixed complex Hilbert space. Let $\mathcal{B}(\mathcal{H})$ be the space of all bounded linear operators on $\mathcal{H}$, and let $\mathcal{A}$ to be the free semigroup $\mathcal{B}(\mathcal{H})$-algebra on $x,$ i.e., $\mathcal{A}=\mathcal{B}(\mathcal{H})\langle x\rangle$. An element $p$ of $\mathcal{A}$ takes the form,
@@ 3.2. Left creation operators
$\iota^{*}L_{i}\iota e_{w}=\iota^{*}L_{i}e_{w}=\begin{cases}e_{x_{j}w}&\text{ if }|w|<\lcal{d}\\ 0&\text{ if }|w|=\lcal{d}.\end{cases}$ (3.2)
@@
$\iota^{*}L_{j}^{*}\iota e_{v}=\iota^{*}L_{j}^{*}e_{v}=\begin{cases}e_{u}&\text{ if }v=x_{j}u\\ 0&\text{otherwise}.\end{cases}$ (3.3)
Let AA
$A_{j}=\iota^{*}B_{j}\iota=\iota^{*}\left(L_{j}+L_{j}^{*}\right)\iota.$ (3.4)
From equations (3.2) and (3.3), $A_{j}w=B_{j}w$ for $|w|<\lcal{d}.$ While $A$ depends on $\lcal{d}$, we suppress this dependence in the notation for readability; when $\lcal{d}$ is not clear from the context, we write $A=A^{(\lcal{d})}$.
@@ 4.1.1. Closedness of $\mathcal{C}_{d}$ in the product ultraweak topology
Equip $\mathcal{A}_{\lcal{d}}$ with a norm in the following way. For $p=\sum\limits_{w\in\langle x\rangle_{\lcal{d}}}P_{w}w$,
$\|p\|\ =\ \sum\limits_{w\in\langle x\rangle_{\lcal{d}}}\|P_{w}\|,$
where $\|T\|$ denotes the operator norm for $T\in\mathcal{B}(\mathcal{H}).$
@@
Fix $t>0.$ Let $(p_{\alpha})_{\alpha}$ be a net in $\mathcal{C}_{d,t}$ that converges to $p\in\mathcal{A}_{2d}.$ Our aim is to show that $p\in\mathcal{C}_{d,t}.$ Note that for any $\alpha$,
$\|p_{\alpha}(A)\|\ \leq\ \|p_{\alpha}\|\,\|A\|\ \leq\ t\|A\|.$
Thus,
$\sup\limits_{\alpha}\|p_{\alpha}(A)\|\ <\ \infty.$
By Proposition 3.3,
$\sup\limits_{\alpha}\|G_{\alpha}\|\ <\ \infty$
for any $G_{\alpha}\in\Gamma_{p_{\alpha}}$. Thus, by the Banach–Alaoglu theorem, there is a subnet of operators $(G_{\beta})_{\beta}$ that converges to some operator $G\in\mathcal{B}(\mathcal{H})^{N(d)\times N(d)}$ in the ultraweak topology. Since the convergence in the ultraweak operator topology is stronger than convergence in the weak operator topology, $G_{\beta}\longrightarrow G$ in the WOT. Thus, $G\succeq 0.$
@@ 7.2. A modified GNS construction
$\mathcal{M}_{y}\ =\ {\rm span}\{\Phi(w)\xi:w\in\left((\mathbb{F}_{{\tt{g}}})_{d-1}\cup y^{-1}(\mathbb{F}_{{\tt{g}}})_{d-1}\right),\,\xi\in\mathcal{H}\}\ \subseteq\ \mathcal{E}.$

### v3

@@ 1.1. Notation
Let $\mathcal{H}$ be a fixed complex Hilbert space. Let $\mathcal{B}(\mathcal{H})$ be the space of all bounded linear operators on $\mathcal{H}$, and let $\mathcal{A}$ to be the free semigroup $\mathcal{B}(\mathcal{H})$-algebra on $x,$ i.e., $\mathcal{A}=\mathcal{B}(\mathcal{H})\langle x\rangle=\mathcal{B}(\mathcal{H})\otimes\mathbb{C}\langle x\rangle$. An element $p$ of $\mathcal{A}$ takes the form,
@@ 3.2. Left creation operators
$\iota^{*}L_{i}\iota e_{w}=\iota^{*}L_{i}e_{w}=\begin{cases}e_{x_{i}w}&\text{ if }|w|<\lcal{d}\\ 0&\text{ if }|w|=\lcal{d}.\end{cases}$ (3.2)
@@
$\iota^{*}L_{i}^{*}\iota e_{v}=\iota^{*}L_{i}^{*}e_{v}=\begin{cases}e_{u}&\text{ if }v=x_{i}u\\ 0&\text{otherwise}.\end{cases}$ (3.3)
Let $A$ be the $\tt{g}$}-tuple of operators defined as
$A_{i}=\iota^{*}B_{i}\iota=\iota^{*}\left(L_{i}+L_{i}^{*}\right)\iota.$ (3.4)
From equations (3.2) and (3.3), $A_{i}w=B_{i}w$ for $|w|<\lcal{d}.$ While $A$ depends on $\lcal{d}$, we suppress this dependence in the notation for readability; when $\lcal{d}$ is not clear from the context, we write $A=A^{(\lcal{d})}$.
@@ 4.1.1. Closedness of $\mathcal{C}_{d}$ in the product ultraweak topology
Let $\mathcal{T}(\mathcal{H})$ be the space of trace-class operators on $\mathcal{H}.$ The space $\mathcal{T}(\mathcal{H})^{\langle x\rangle_{\lcal{d}}}$ with the norm
$\|(T_{1},\dots,T_{N(\lcal{d})})\|_{1}\ =\ \sum\limits_{i=1}^{N(\lcal{d})}\|T_{i}\|_{1},$
is a Banach space, where $\|\cdot\|_{1}$ denotes the trace-class norm of $\mathcal{T}(\mathcal{H})$. The dual of $\mathcal{T}(\mathcal{H})^{\langle x\rangle_{\lcal{d}}}$ is $\mathcal{B}(\mathcal{H})^{\langle x\rangle_{\lcal{d}}}$ with the induced norm
$\|(T_{1},\dots,T_{N(\lcal{d})})\|\ =\ \max\limits_{1\leq i\leq N(\lcal{d})}\,\|T_{i}\|,$
where $\|\cdot\|$ denotes the operator norm of $\mathcal{B}(\mathcal{H}).$ Via the canonical identification, we can endow $\mathcal{A}_{\lcal{d}}$ with a norm. For $p=\sum\limits_{w\in\langle x\rangle_{\lcal{d}}}P_{w}w$, we define
$\|p\|\ :=\ \max\limits_{w\in\langle x\rangle_{\lcal{d}}}\,\|P_{w}\|.$
@@
Fix $t>0.$ Let $(p_{\alpha})_{\alpha}$ be a net in $\mathcal{C}_{d,t}$ that converges to $p\in\mathcal{A}_{2d}.$ Our aim is to show that $p\in\mathcal{C}_{d,t}.$
For each $\alpha$,
$\|p_{\alpha}(A)\|\ \leq\ \|p_{\alpha}\|\,\sum\limits_{w\in\mathcal{A}_{2d}}\|A^{w}\|\ \leq\ t\,\sum\limits_{w\in\mathcal{A}_{2d}}\|A^{w}\|.$
Hence,
$\sup\limits_{\alpha}\|p_{\alpha}(A)\|\ <\ \infty.$ (4.1)
Choose $G_{\alpha}\in\Gamma_{p_{\alpha}}.$ Combining equation (4.1) and Proposition 3.3 gives,
$\sup\limits_{\alpha}\|G_{\alpha}\|\ <\ \infty.$
Thus, by the Banach–Alaoglu theorem, there is a subnet of operators $(G_{\beta})_{\beta}$ that converges to some operator $G\in\mathcal{B}(\mathcal{H})^{N(d)\times N(d)}$ in the ultraweak topology. Since the convergence in the ultraweak operator topology is stronger than convergence in the weak operator topology, $G_{\beta}\longrightarrow G$ in the WOT. Thus, $G\succeq 0.$
@@ 7.2. A modified GNS construction
$\mathcal{M}_{y}\ =\ {\rm span}\{\Phi(w)\xi:w\in\left((\mathbb{F}_{{\tt{g}}})_{d-1}\cup y(\mathbb{F}_{{\tt{g}}})_{d-1}\right),\,\xi\in\mathcal{H}\}\ \subseteq\ \mathcal{E}.$
@@ 8. Concluding remark and a problem
Declarations
Competing interests
The authors have no relevant financial or non-financial interests to disclose.
Data availability
This article is theoretical in nature; no datasets were generated or analyzed. All results and proofs are contained within the article.